A total of $16,000 is invested in two individual retire-ment accounts paying 5% and 6% annual interest. The combined annual interest earned is $860. How much of the $16,000 is invested in each account? (Hint: you will need to use the formula I = prt; design one equation for the amount invested in each fund and another for the interest earned.
Answer by mananth(16946)

Part I 5.00% per annum ------------- Amount invested =x
Part II 6.00% per annum ------------ Amount invested = y
16000
Interest----- 860.00

Part I 5.00% per annum ---x
Part II 6.00% per annum ---y
Total investment
x + 1 y= 16000 -------------1
Interest on both investments
5.00% x + 6.00% y= 860
Multiply by 100
5 x + 6 y= 86000.00 --------2
Multiply (1) by -5
we get
-5 x -5 y= -80000.00
Add this to (2)
0 x 1 y= 6000
divide by 1
y = 6000
Part I 5.00% $ 10000
Part II 6.00% $ 6000

CHECK
10000 --------- 5.00% ------- 500.00
6000 ------------- 6.00% ------- 360.00
Total -------------------- 860.00

m.ananth@hotmail.ca